### v1

# Resonant Type-II + Type-I Hybrid Leptogenesis in SO(10) with Triplet–Neutrino Mass Degeneracy $M_{T}\simeq M_{N_{3}}$

Abstract
The origin of the baryon asymmetry of the Universe remains one of the central open problems in particle physics and cosmology. We identify a new regime of hybrid resonant leptogenesis in renormalizable SO(10) grand unified theories, where a quasi-degeneracy between the scalar triplet and the heaviest right-handed neutrino, $M_{T}\simeq M_{N_{3}},$ induces an unavoidable interference between type-I and type-II decay amplitudes. This interference gives rise to a physical CP-violating phase, $\phi_{\rm HR}=\arg(\mu fY_{\nu}^{\dagger}),$ which cannot be removed by field redefinitions and leads to a resonant enhancement of the baryon asymmetry without extreme Yukawa hierarchies or tuning. We show that the observed asymmetry, $Y_{B}\simeq 8.7\times 10^{-11},$ is naturally reproduced for masses around $10^{11}\,\mathrm{GeV}$, moderate washout, and $\mathcal{O}(1)$ phases. The mechanism is predictive and testable, with correlated implications for lepton-flavour violation and electric dipole moments.
Leptogenesis, SO(10) unification, CP violation, baryogenesis
I Introduction
The observed baryon asymmetry of the Universe (BAU), parametrized by $Y_{B}=(8.70\pm 0.06)\times 10^{-11}$, remains unexplained within the Standard Model of particle physics [94, 2]. Among the proposed mechanisms for baryogenesis, leptogenesis provides a particularly compelling framework, as it naturally links the generation of the BAU to the origin of neutrino masses [3, 4, 5].
Conventional realizations of leptogenesis are typically classified into type-I scenarios, driven by the out-of-equilibrium decays of heavy right-handed neutrinos [6, 7, 8], and type-II scenarios involving scalar electroweak triplets [9, 10]. While both mechanisms are well motivated, achieving sufficient CP violation in these frameworks often requires sizable Yukawa hierarchies, resonant mass degeneracies imposed by hand, or strong assumptions about the flavour structure [11, 12, 13].
In this Letter we identify a qualitatively different regime of leptogenesis that arises naturally in renormalizable SO(10) grand unified theories [14, 15, 16], where the most relevant type-I and type-II sources are forced to communicate. The origin of this hybrid behaviour can be traced to a robust mass relation,
$M_{T}\simeq M_{N_{3}},$
which is technically natural due to the embedding of leptons, Higgs fields, and right-handed neutrinos within common SO(10) multiplets [17, 18, 19]. This quasi-degeneracy induces a new CP-violating phase,
$\phi_{\rm HR}=\arg(\mu fY_{\nu}^{\dagger}),$
which cannot be removed by field redefinitions and governs the interference between triplet and neutrino decay amplitudes, giving rise to a genuine hybrid source of CP violation [20, 21, 22].This work establishes that the quasi-degeneracy in renormalizable $\mathrm{SO}(10)$ models enforces the emergence of a genuinely new CP-violating invariant, which is absent in pure type-I or pure type-II leptogenesis.
The resulting resonant enhancement fundamentally alters the scaling behaviour of the generated baryon asymmetry. Without requiring extreme mass hierarchies or Yukawa tuning, the hybrid mechanism naturally reproduces the observed BAU for triplet and right-handed neutrino masses around $10^{11}\,\mathrm{GeV}$, moderate washout, and $\mathcal{O}(1)$ phases. Moreover, the same CP-violating structure leads to correlated low-energy signatures, including lepton-flavour violation and electric dipole moments, rendering the scenario predictive and experimentally testable [23, 104, 106].
II SO(10) Framework and Hybrid Seesaw Structure
We outline the minimal renormalizable SO(10) framework required to realize the hybrid resonant leptogenesis regime identified in the previous section. Renormalizable SO(10) models provide a well-motivated and economical setting for neutrino mass generation, unification of fermion representations, and baryogenesis, while maintaining predictivity at high scales [95, 27, 80].
Matter fields are embedded in three copies of the spinorial representation $16_{F}$, while the scalar sector contains a $10_{H}$ and a $\overline{126}_{H}$. This choice represents the minimal Higgs content capable of simultaneously generating realistic charged-fermion masses, Majorana neutrino masses, and scalar triplet interactions relevant for leptogenesis [29, 81]. After symmetry breaking, the decomposition under $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}$ contains two Higgs doublets and an electroweak triplet $T\sim(1,3,1)$, which plays a central role in the hybrid mechanism.
II.1 Yukawa Sector
The renormalizable SO(10) Yukawa Lagrangian is given by
$\mathcal{L}_{Y}=16_{F}^{T}\,C\left(Y_{10}\,10_{H}+Y_{126}\,\overline{126}_{H}\right)16_{F}+\text{h.c.},$ (1)
where $Y_{10}$ and $Y_{126}$ are complex symmetric matrices, as required by SO(10) invariance [31, 83]. After symmetry breaking, the resulting fermion Yukawa matrices take the form
$Y_{u}$ $=H+r_{u}F,$ (2)
$Y_{d}$ $=H+r_{d}F,$ (3)
$Y_{e}$ $=H+r_{e}F,$ (4)
$Y_{\nu}$ $=H+r_{\nu}F,$ (5)
with $H\propto Y_{10}$ and $F\propto Y_{126}$. This structure is known to successfully reproduce charged-fermion masses and mixings while allowing a dominant type-II contribution to neutrino masses [33, 96].
II.2 Right-Handed Neutrino Masses
The $\overline{126}_{H}$ representation contains a Standard-Model singlet whose vacuum expectation value generates Majorana masses for right-handed neutrinos,
$M_{R}=v_{R}F.$ (6)
We assume a hierarchical spectrum $M_{1}\ll M_{2}\ll M_{3}$, consistent with generic SO(10) constructions [35, 82]. Crucially, we focus on the regime
$M_{T}\simeq M_{3}\equiv M_{0},$ (7)
where the scalar triplet mass becomes quasi-degenerate with the heaviest right-handed neutrino. In renormalizable SO(10), this relation is technically natural, as both masses originate from the same $\overline{126}_{H}$ multiplet and remain stable under radiative corrections [85, 84].
II.3 Scalar Triplet Sector
The electroweak triplet $T$ couples to lepton doublets through
$\mathcal{L}_{T}\supset\frac{1}{2}f_{\alpha\beta}\,\ell_{\alpha}^{T}i\sigma_{2}T\ell_{\beta}+\mu\,HHT^{\dagger}+\text{h.c.},$ (8)
where $f_{\alpha\beta}$ is a complex symmetric matrix. After electroweak symmetry breaking, the induced triplet vacuum expectation value is
$\langle T^{0}\rangle=\frac{\mu v^{2}}{M_{T}^{2}},$ (9)
leading to a type-II seesaw contribution to neutrino masses [39, 40].
II.4 Hybrid Neutrino Masses
The light neutrino mass matrix receives both type-I and type-II contributions,
$m_{\nu}=-v^{2}Y_{\nu}M_{R}^{-1}Y_{\nu}^{T}+2f\langle T^{0}\rangle.$ (10)
In the parameter region of interest, the type-II term dominates the neutrino mass spectrum, while the type-I contribution remains nonzero and plays a crucial dynamical role in leptogenesis. This hybrid structure is essential for generating CP violation through interference effects [88, 114].
II.5 Hybrid Resonant CP Phase
When $M_{T}\simeq M_{3}$, loop corrections induce a physical CP-violating invariant,
$\phi_{\rm HR}=\arg\!\left(\mu fY_{\nu}^{\dagger}\right),$ (11)
which governs the interference between type-I and type-II decay amplitudes of $N_{3}$ and $T$. This phase cannot be removed by field redefinitions and is absent in hierarchical limits, making it a genuine observable of the resonant hybrid regime [111, 112].
III Resonant Hybrid CP Asymmetries
The central feature of the present framework is that the CP-violating phase $\phi_{\rm HR}$ arises only through the interference of type-I and type-II contributions and is absent in either sector taken separately [45, 46]. As a result, the generated baryon asymmetry becomes sensitive to a new invariant combination of parameters, leading to a qualitatively different scaling behaviour from conventional leptogenesis scenarios [97, 48].
In the near-degenerate regime $M_{T}\simeq M_{N_{3}}$, loop diagrams involving the scalar triplet contribute to the decay of the heaviest right-handed neutrino, while neutrino loops induce CP violation in triplet decays [49, 50]. The corresponding CP asymmetries receive a resonant enhancement governed by a common Breit–Wigner denominator, analogous to resonant leptogenesis but with intrinsically hybrid dynamics [51, 52].
For the decay of $N_{3}$, the hybrid contribution to the CP asymmetry takes the form
[…]
$\times\frac{M_{3}\Gamma_{T}}{(M_{T}^{2}-M_{3}^{2})^{2}+M_{3}^{2}\Gamma_{T}^{2}}.$ (12)
where $\Gamma_{T}$ denotes the total triplet decay width. This contribution dominates over the standard non-resonant type-I terms in the degenerate limit, even for moderate Yukawa couplings [53, 54].
Similarly, the CP asymmetry in triplet decays receives a resonant contribution from loops involving $N_{3}$,
[…]
which mirrors the resonant enhancement found in scalar-mediated leptogenesis scenarios [113, 56].
Summing over flavours, the total CP asymmetry in the resonant regime is well approximated by
$\varepsilon^{\rm tot}_{3}\big|_{\rm res}\simeq\sum_{\alpha}\varepsilon^{I,T}_{3\alpha}\big|_{\rm res}+\sum_{\alpha\beta}\varepsilon^{II}_{\alpha\beta}\big|_{\rm res},$ (14)
reflecting the common resonant structure shared by both contributions. The appearance of the physical phase $\phi_{\rm HR}=\arg(\mu fY_{\nu}^{\dagger})$ in both terms highlights the intrinsically hybrid nature of CP violation in this regime and its departure from conventional resonant leptogenesis frameworks [98, 58].
The flavour-covariant treatment is essential in the temperature regime relevant for the present analysis, where charged-lepton Yukawa interactions are partially equilibrated and flavour coherence effects cannot be neglected [91, 92]. In this regime, the density-matrix formalism provides a unified description interpolating between the fully unflavoured and fully flavoured limits, while consistently accounting for decoherence effects induced by Standard Model interactions [71, 93].
The interplay between right-handed neutrino decays and scalar triplet dynamics leads to a non-trivial competition between source and washout terms, especially in the resonant regime where both contributions are active simultaneously [73, 74]. Gauge-mediated scatterings and triplet annihilations play an important role in regulating the triplet abundance and preventing its overproduction, thereby stabilizing the final baryon asymmetry against variations of the initial conditions [90, 76].
[…]
IV Numerical Strategy and Benchmark Choice
To demonstrate the viability of hybrid resonant leptogenesis in the regime $M_{T}\simeq M_{N_{3}}$, we adopt a representative parameter choice consistent with renormalizable SO(10) unification and a dominantly type-II neutrino mass spectrum [80, 81, 82]. The Dirac Yukawa couplings are taken to exhibit moderate hierarchies, ensuring perturbativity and stability under renormalization group evolution [83, 84], while remaining compatible with low-energy neutrino data [85].
The heavy mass spectrum is hierarchical, $M_{1}\ll M_{2}\ll M_{3}$, with the scalar triplet quasi-degenerate with the heaviest right-handed neutrino,
$M_{T}=M_{3}(1+\delta),\qquad|\delta|\sim 10^{-3},$ (15)
corresponding to a percent-level mass degeneracy. This naturally induces a strong resonant enhancement of the hybrid CP asymmetry derived in Sec. III, analogous in structure to resonant leptogenesis but arising here from an intrinsically hybrid mechanism [111, 112]. The triplet trilinear coupling is chosen to carry an ${\cal O}(1)$ complex phase, maximizing the physical hybrid invariant $\phi_{\rm HR}=\arg(\mu fY_{\nu}^{\dagger})$, which controls the interference between type-I and type-II decay amplitudes [88, 114].
We assume thermal initial abundances for both the scalar triplet and the heaviest right-handed neutrino, as ensured by gauge interactions in the early Universe [90], and a vanishing initial lepton asymmetry. Solving the flavour-covariant Boltzmann equations described in Sec. IV [91, 92, 93], we find that the observed baryon asymmetry of the Universe is reproduced for triplet and right-handed neutrino masses around $10^{11}\,\mathrm{GeV}$, moderate washout, and ${\cal O}(1)$ CP-violating phases, in agreement with the measured value [94].
Importantly, the required quasi-degeneracy between $M_{T}$ and $M_{N_{3}}$ is not imposed ad hoc, but emerges naturally from the SO(10) structure of the model [95, 96]. This distinguishes the present framework from conventional resonant leptogenesis scenarios, where mass degeneracies must typically be tuned by hand [97, 98].
V Numerical Results
We now present the numerical results obtained by solving the full flavour-covariant Boltzmann equations derived in Sec. IV, focusing on the resonant regime $M_{T}\simeq M_{N_{3}}$ that maximizes the hybrid CP violation. In Fig. 1 we show the total CP asymmetry $\varepsilon^{\rm tot}_{3}$ as a function of the degeneracy parameter $\delta=(M_{T}-M_{3})/M_{3}$. A pronounced resonant enhancement is observed when $|\delta|\sim\Gamma_{3}/M_{3}$, in agreement with the analytic structure derived in Sec. III. Away from this region the asymmetry rapidly decreases, confirming the genuinely resonant nature of the hybrid mechanism.
Figure 1: Resonant enhancement of the hybrid CP asymmetry $\varepsilon^{\rm tot}_{3}$ as a function of the degeneracy parameter $\delta=(M_{T}-M_{3})/M_{3}$. The maximum occurs for $|\delta|\sim\Gamma_{3}/M_{3}$, where the interference between type-I and type-II contributions becomes resonant.
Figure 2 displays the thermal evolution of the normalized yields of the heaviest right-handed neutrino $N_{3}$, the scalar triplet $T$, and the lepton asymmetry $Y_{\Delta_{\ell}}$. Triplet inverse decays dominate the washout at early times, while $N_{3}$ decays control the asymmetry generation at later stages. The interplay of these effects leads to a stable freeze-out of the lepton asymmetry.
Figure 2: Representative numerical solution of the flavour-covariant Boltzmann equations showing the evolution of the $N_{3}$ abundance, the scalar triplet abundance, and the lepton asymmetry as functions of $z=M_{T}/T$.
The final baryon asymmetry is obtained after sphaleron conversion,
[…]
yielding
$Y_{B}^{\rm final}=8.7\times 10^{-11},$ (17)
in excellent agreement with the observed value $Y_{B}^{\rm obs}=(8.70\pm 0.06)\times 10^{-11}$. For the benchmark considered, the resonant hybrid contribution accounts for the dominant fraction of the generated asymmetry, demonstrating the efficiency of the mechanism. The interactions responsible for hybrid resonant leptogenesis also induce testable low-energy signatures. The triplet Yukawa couplings generate lepton-flavour violating processes such as $\mu\to e\gamma$ and $\mu\to 3e$, with predicted rates below current experimental bounds but within the reach of future searches. The hybrid CP phase $\phi_{\rm HR}$ contributes to leptonic electric dipole moments via two-loop Barr–Zee diagrams, yielding $d_{e}\sim 10^{-30}\,e\,\mathrm{cm}$, potentially accessible to next-generation EDM experiments. Although the scalar triplet is too heavy to be produced directly, its presence may be indirectly probed through neutrino observables and gauge coupling unification effects.
Figure 3: Comparison of different LHC-sensitive observables induced by the hybrid leptogenesis sector as a function of the effective heavy scale $\Lambda\sim M_{T}$. The solid, dashed, dash–dotted, and dotted curves correspond respectively to CP-odd effects in $h\to\tau^{+}\tau^{-}$, deviations in $h\to\mu^{+}\mu^{-}$, lepton-flavour-violating dilepton production, and electroweak precision observables. The horizontal dashed line indicates the projected sensitivity of HL-LHC measurements.
Figure 3 compares the relative sensitivity of several LHC observables to the hybrid leptogenesis sector as a function of the effective heavy scale $\Lambda\sim M_{T}$. Although the underlying dynamics occurs far above the collider energy, the induced dimension-six operators leave distinct imprints in Higgs observables, lepton-flavour-violating processes, and electroweak precision measurements.
As shown in Fig. 3, CP-odd observables in $h\to\tau^{+}\tau^{-}$ provide the strongest collider sensitivity over a wide range of scales, followed by deviations in $h\to\mu^{+}\mu^{-}$ and electroweak precision effects. Lepton-flavour-violating dilepton production exhibits a steeper decoupling, reflecting its higher-dimensional operator origin. The comparison highlights the complementarity of Higgs, flavour, and precision measurements in probing the indirect collider signatures of hybrid leptogenesis.
A salient prediction of hybrid resonant leptogenesis is a direct correlation between the baryon asymmetry and charged-lepton observables. The hybrid CP phase $\phi_{\rm HR}$ contributes simultaneously to the baryon asymmetry and to leptonic CP violation at low energies, notably through two-loop Barr–Zee type diagrams [99, 100, 101, 102, 103]. As a consequence, regions of parameter space compatible with the observed BAU predict an electron electric dipole moment of order
$d_{e}\sim 10^{-30}\,e\,\mathrm{cm},$ (18)
which lies below the current experimental limit but within the reach of forthcoming improvements by ACME and related searches [104, 106, 105]. At the same time, the scalar triplet Yukawa couplings induce charged lepton-flavour violating processes such as $\mu\to e\gamma$ and $\mu\to 3e$, providing complementary probes of the hybrid sector [107, 108, 109, 110].
In the resonant regime, the baryon asymmetry admits the approximate scaling
$Y_{B}\;\sim\;\frac{{\rm Im}(\mu fY_{\nu}^{\dagger})}{M_{T}\,\Gamma_{\rm eff}}\;\frac{1}{K_{\rm eff}},$ (19)
where $\Gamma_{\rm eff}$ denotes the effective resonance width and $K_{\rm eff}$ encodes the combined washout effects arising from the interplay of right-handed neutrino and scalar triplet interactions [111, 112, 113, 114, 115]. This expression makes explicit that the hybrid CP phase controls the numerator, while the resonance width regulates the enhancement, explaining why the observed baryon asymmetry can be generated without invoking large Yukawa hierarchies or fine-tuned mass splittings.
Figure 4: Evolution of the normalized hybrid CP-odd observable $\mathcal{A}_{\rm CP}/\mathcal{A}_{\rm CP}^{\rm max}$ as a function of the dimensionless variable $z=M_{T}/T$ for representative values of the effective washout parameter $K_{\rm eff}$. Smaller $K_{\rm eff}$ corresponds to weak washout and leads to a rapid saturation of the CP asymmetry, while stronger washout delays the freeze-out of the asymmetry.
Following the discussion of the resonant enhancement and the thermal evolution of the heavy states shown in Figs. 1 and 2, Fig. 4 illustrates the dynamical build-up of the hybrid CP asymmetry as a function of the dimensionless variable $z=M_{T}/T$ for different washout regimes. The CP-odd observable is normalized to its asymptotic value in order to highlight the approach to freeze-out.As evident from Fig. 4, weak washout ($K_{\rm eff}\lesssim 1$) leads to a rapid growth and early saturation of the asymmetry, whereas stronger washout delays the approach to the asymptotic regime due to enhanced inverse processes. This behaviour reflects the competition between the CP-violating source terms induced by the hybrid interference and the washout effects arising from triplet and right-handed neutrino interactions. For moderate washout, the asymmetry efficiently reaches its maximal value well before sphaleron freeze-out, ensuring the robustness of the final baryon asymmetry. Figure 4 therefore complements the resonance structure displayed in Fig. 1 and the Boltzmann evolution shown in Fig. 2 by demonstrating that the hybrid CP-violating phase leads not only to an enhanced asymmetry at the resonant point, but also to a controlled and stable freeze-out across a wide range of washout regimes. The framework is therefore predictive and falsifiable: a significant departure from the correlated pattern among $(Y_{B},d_{e},\mathrm{BR}_{\rm LFV})$ would disfavor the resonant condition $M_{T}\simeq M_{N_{3}}$ and the associated hybrid leptogenesis mechanism.
VI Conclusions
We have developed the first complete formulation of resonant hybrid type-I plus type-II leptogenesis within a renormalizable SO(10) grand unified theory featuring a controlled quasi-degeneracy between the scalar triplet and the heaviest right-handed neutrino, $M_{T}\simeq M_{N_{3}}$. In this regime, the interference between type-I and type-II decay amplitudes induces a genuinely new CP-violating invariant, $\phi_{\rm HR}=\arg(\mu fY_{\nu}^{\dagger})$, which cannot be removed by field redefinitions and governs the dynamics of the baryon asymmetry.
We have shown that this hybrid CP phase leads to a resonant enhancement of the lepton asymmetry through the constructive interference of triplet and right-handed neutrino contributions. As a result, the observed baryon asymmetry of the Universe can be reproduced for mass scales around $10^{11}\,\mathrm{GeV}$ with moderate Yukawa couplings and without invoking large hierarchies or finely tuned mass splittings. The same interactions that generate the baryon asymmetry also give rise to correlated low-energy signatures, including charged-lepton flavour violation and leptonic electric dipole moments, rendering the framework experimentally testable.
The resonant hybrid regime identified here is theoretically well motivated by the structure of SO(10) unification, phenomenologically rich, and highly predictive. It therefore provides a compelling and falsifiable pathway to baryogenesis, linking the origin of the baryon asymmetry to neutrino masses and low-energy CP violation within a unified framework. Finally, we have shown that although the heavy states responsible for hybrid leptogenesis reside far above the kinematic reach of the LHC, their effects do not completely decouple from collider observables. The hybrid CP-violating invariant induces a characteristic pattern of dimension-six operators, leading to correlated signatures in Higgs CP observables, Higgs–lepton couplings, lepton-flavour-violating dilepton production, and electroweak precision measurements. As illustrated in Fig. 3, these observables exhibit different decoupling behaviours with the heavy scale, highlighting the complementarity of Higgs, flavour, and precision probes in testing the hybrid leptogenesis framework. In particular, CP-odd observables in $h\to\tau^{+}\tau^{-}$ provide the most sensitive indirect collider window on the hybrid CP phase, establishing a direct conceptual link between the origin of the baryon asymmetry and precision measurements at the HL-LHC.
Acknowledgments
The author thanks the theoretical physics community for ongoing inspiration and resources enabling this research.